Minimum Wage. Determine the job loss.
D: P = 1,640 - 2Q
S: P = 240 + 2Q
MW: P = 1,260
1
Expert's answer
2016-05-22T12:42:02-0400
D: P = 1,640 - 2Q
S: P = 240 + 2Q
In equilibrium Qd = Qs and Pd = Ps, so:
1,640 - 2Q = 240 + 2Q,
4Q = 1,400,
Qe = 350 persons.
Pe = 240 + 2*350 = 940 persons.
If the minimum wage is P = 1,260, then quantity demanded is:
1,260 = 1,640 - 2Q,
2Q = 380,
Q = 190 persons.
The quantity supplied is:
1,260 = 240 + 2Q,
2Q = 1,020,
Q = 510 persons.
So, the job loss is Qe - Qd = 350 - 190 = 160 persons.